(fog)(x):f(x)=8x+13,g(x)=2x-1
Answer by stanbon(75887)

(fog)(x):f(x)=8x+13,g(x)=2x-1
--------
fog(x) = f[g(x)] = f[2x-1] = 8(2x-1)+13 = 16x+5
=========================
Cheers,
Stan H.
=============